OpenStudy (drmoss):

Dante is making a necklace with 18 rows of tiny beads in which the number of beads per row is given by the series 3 + 10 + 17 + 24 + ... a. Use summation notation to write the series, give i.the lower limit of the sum ii.the upper limit of the sum iii.the explicit formula of the sum b. Find the total number of beads in the necklace. Explain your method for finding the total number of beads.

OpenStudy (drmoss):

I only need help with part a iii

OpenStudy (drmoss):

a(18)=3+(3-1)7 would that be right?

OpenStudy (drmoss):

@jhonyy9 @Anaise @alexalc @brian.edu @DiamondAnon2 @sammixboo

OpenStudy (drmoss):

lol calling everyone!

jhonyy9 (jhonyy9):

do you see how many is the difference between two terms ?

OpenStudy (drmoss):

yeah can u check my work.

jhonyy9 (jhonyy9):

where is ?

OpenStudy (drmoss):

first comment 2nd comment

jhonyy9 (jhonyy9):

ok a_1 = ? what will be the right formula of this series ?

jhonyy9 (jhonyy9):

and so a_n = ?

jhonyy9 (jhonyy9):

a_1 = 3 a_2 = a_1 +7 = 3+7 = 10 a_3 = a_2 +7 = 10 +7 = 17 ----- a_18 = ?

jhonyy9 (jhonyy9):

what you wrote above for a_18 not can being right because from there result 17 and thatn you check it the term a_3 is equal 17 so a_18 not can being 17 too

OpenStudy (drmoss):

i dont get it

OpenStudy (drmoss):

can you tell me what the equation is?

jhonyy9 (jhonyy9):

sorry

OpenStudy (drmoss):

?

OpenStudy (drmoss):

ohh i get it i think

jhonyy9 (jhonyy9):

a_1 = 3 a_2 = 10 = a_1 +7 a_3 = 17 = a_2 +7

OpenStudy (drmoss):

a(18)=a(17)+7

jhonyy9 (jhonyy9):

yes this i ve wrote above too for term a_18 than you check it there

OpenStudy (drmoss):

ok ty

jhonyy9 (jhonyy9):

@DiamondAnon2 what is your idea please ? - write it here - thank you

OpenStudy (diamondanon2):

Oh I did not want to interrupt, but I have another question.

jhonyy9 (jhonyy9):

bc. i saw you started writing anything - yes ?

jhonyy9 (jhonyy9):

@DiamondAnon2 please

OpenStudy (diamondanon2):

I just have another question. But i do not want to stop you from helping moss.

jhonyy9 (jhonyy9):

ok. post it there and tag me there please - and i will try helping you too

jhonyy9 (jhonyy9):

Moss ok ?